OpenStudy (ashy98):

9 subtracted from x is equal to x divided by 4??? What is x???

OpenStudy (anonymous):

9 subtracted from x is x-9. x divided by 4 is x/4. x-9=x/4.now do you know what to do first?

OpenStudy (anonymous):

\[x-9=x \div4\] \[-9=(x/4)-x\] \[-36=x-4x\] \[-36=3x\] \[-36/3=x\] \[x=12\]

OpenStudy (anonymous):

do you understand how i did that? @Ashy98

OpenStudy (dan815):